# A Simple and Explainable Model for Park-and-Ride Car Park Occupancy Prediction

### v1

@@ 4.3.1 Fitting the TN Model
For the basic TN model, the objective is to estimate the parameters $\boldsymbol{\theta}=(\mu_{a},\mu_{d},\sigma_{a},\sigma_{d})$ from a dataset consisting of $N$ daily occupancy profiles, $\mathcal{D}~{}=~{}\{o_{1:T}\}_{n=1}^{N}$, recorded at $T$ discrete timestamps [^2] In our case, $T=48$ since the dataset has a 30-minute time resolution.. To ensure consistency, all occupancy profiles are normalized such that $\sum_{i=1}^{T}o_{i}=1$.
@@
The optimization can be carried by minimizing the sum of squares loss
@@ 4.3.2 Fitting the TNL Model
For the TNL model with occupancy limit, the set of fitted parameters $\boldsymbol{\theta}$ also includes a set of threshold parameters $\{\tau_{i}\}_{i=1}^{N}$. Thus, we allow $\tau$ to take a different value for each training day $i$ while keeping the other parameters fixed. This approach ensures that the parameters of the arrival distribution remain independent of the exact moment a parking lot reaches capacity. As a result, the model can not only predict when the capacity limit will be reached but also estimate the number of cars that were unable to find a free parking space. Note, however, that this approach requires a different normalization of the data. In this case, all occupancy profiles are scaled such that $\max(o_{1:T})=1$. This normalization is essential for identifying the moments when the parking lots reach full capacity.
@@ 4.4.1 Average Activity Profile
A straightforward method for predicting patterns influenced by circadian rhythms is the use of average activity profiles, as employed in [11]. We adapt this approach by computing the average over normalized, aggregated data, distinguishing between three groups: weekdays (Monday to Thursday), Fridays, and weekends (Saturday and Sunday).
@@ 4.4.2 Linear Regression
The regression model optimizes the following objective:
@@ 5.1 Model Fit Quality
We begin by assessing how well the model represents both the training and test data. For this, we use the mean relative proportional error, computed relative to the maximum occupancy of a parking lot. For clarity, we group the test data into three categories based on observed activity patterns (as illustrated in Figure 3): weekdays (Monday to Thursday), Fridays, and weekends (Saturday and Sunday).
@@ 5.1.1 Example Fits on Training Data
Figure 7: Example of the TN model fitted to training data for the Vilanova car park. The corresponding parameters of the arrival and departure TN distributions are $\mu_{a}=~{}$06:55h, $\mu_{d}=~{}$18:40 for weekdays, $\mu_{a}=~{}$07:02h, $\mu_{d}=~{}$17:26h for Fridays and $\mu_{a}=~{}$07:49h weekends. Note that for the TN model the data is normalized by the area under the curve.
@@
Figure 8: Example of the TNL model fitted to training data for the Quatre Camins car park. The corresponding parameters of the arrival and departure distributions are $\mu_{arr}=~{}$07:31h, $\mu_{dep}=~{}$19:24h for weekdays, $\mu_{arr}=~{}$07:43h, $\mu_{dep}=~{}$18:29h for Fridays and $\mu_{arr}=~{}$10:10h and for weekends. Dashed lines indicate stdv of the model curve. The car park usually reaches its capacity limit from Monday to Friday between 8 and 8:30. Note that for the TNL model the data is normalized by the max value.
@@
We first illustrate model fits on the training data using two example car parks. Figure 7 shows the extracted TN model curves for the Vilanova car park, while Figure 8 shows the TNL model curves, incorporating an occupancy limit, for the Quatre Camins car park. The latter reaches maximum capacity on most weekdays and Fridays (since for the weekend profile this does not happen the basic TN model is used there).
@@
The dark grey lines represent actual occupancy profiles from the training data. The best-fitting model curves are shown as solid red lines, while blue dashed lines indicate the standard deviation of the model fit. All curves are normalized: in the TN model, by the area under the curve, and in the TNL model, by the maximum value. The latter normalization is necessary to estimate excess demand in a given car park. For Figure 8, the model curve is plotted using the average excess across all training curves. We observe the best fits for weekday profiles, while Fridays, especially weekends, exhibit greater variability.
@@
The fitted TN model curves and their corresponding parameters are shown in Figure 9. The figure presents model curves for weekdays (left panel) and Fridays (right panel) across all eight car parks, ordered by the difference between the location parameters $\mu$ of the arrival and departure distributions (represented as circle markers). The interval corresponding to ±$\pm 1$ standard deviation $\sigma$ around $\mu$ is indicated by dashed lines and horizontal markers for both the arrival and departure distributions.
@@
A smaller time difference between arrivals and departures is observed on Fridays, primarily due to earlier departure times. This pattern aligns with the common practice of shorter work schedules on Fridays, often without lunch breaks. The most pronounced example is the Prat del Llobregat car park, where the average departure time shifts from approximately 17:30 to 14:30—a three-hour advancement. We exclude the corresponding curves for weekends, which are characterized by more extreme parameter values and greater variability.
@@
Figure 9: Model curves, with parameters for the TN model, for weekends (left panel) and Fridays (right panel) for the different car parks. The dots indicated the mean $\mu$ and the dashed line the corresponding standard deviation interval $[\mu-\sigma,\mu+\sigma]$ of the underlying TN distributions for arrival (right) and departure (left).
@@ 5.1.2 Evaluating Model Performance on Test Data
We now analyze the model’s performance by evaluating errors in the test data. Examples for two individual stations are given in Appendix A.
@@
We find that, apart from the four stations that show significant variability on Fridays (Prat Del Llobregat, Cerdanyola, Mollet, and Sant Boi, the latter also on weekends), the average error is always below 4%. The high error on weekends for the Sant Boi car park is caused by party-goers visiting nearby nightlife attractions. This results in an additional high-occupancy phase during the early morning hours, which is not well captured by the current models.
@@ 5.2 Prediction
Next, we analyze the predictive quality of our model. More precisely, at a given moment in time, we aim to predict the evolution of occupancy in the subsequent hours of the day. Again, we measure prediction quality using the relative absolute error with respect to the maximum occupancy of a car park and aggregate weekdays into three groups. We also compare the results with average profiles, which represent the typical activity in a given car park for the corresponding types of days in the training data.
@@
The left panels display real data (dash-dotted lines) along with the predicted occupancy, while the right panels show the corresponding prediction errors for both the model curve (in blue) and the average profile (in orange). The average errors are depicted as dashed lines. The dark gray area represents the data used to generate the prediction.
@@
Figure 11: Three examples for predictions in a car park (Vilanova) with the TN model. Left: The grey areas indicate the known activity used to predict the number of parked cars (red dashed-dotted lines) during the remaining hours of the day. Right: evolution of the relative prediction error in %. Blue lines indicate performance of the TN model while orange lines show the comparison with an average day-cycle profile.
Figure 12: Three examples for predictions in a car park which fills up (QuatreCamins) with the TNL model. Left: The grey areas indicate the known activity used to predict the number of parked cars (red dashed-dotted lines) during the remaining hours of the day. Right: evolution of the relative prediction error in %. Blue lines indicate performance of the TNL model while orange lines show the comparison with an average day-cycle profile. Black line gives an estimate of the surplus (how many cars will not or did not fit in the parking).
@@ 5.3 Nowcasting
Table 1: Median nowcasting error in % for the four stations depicted in Figure 14 for weekdays and Fridays comparing the Truncated Normal with Limit (TNL), basic (TN), and Average Profile (Av. Prof.) model with a Linear Regression (L-Reg.). Nowcasting between 7:00 and 15:00 performed in 30-minute intervals.
@@ 6 Conclusion
Additionally, our work differentiates itself from related studies through its specific focus on the modelling approach. Unlike some studies that prioritize prediction using techniques without an underlying behavioural model [3, 4, 5], our approach is more ambitious, aiming to understand the fundamental behavioural patterns of commuters rather than merely forecasting occupancy trends.
@@
Ultimately, we believe this work will contribute to improved urban planning by offering a practical tool for understanding and predicting commuter behaviour in park-and-ride systems. By balancing simplicity, interpretability, and predictive accuracy, our approach can support data-driven decision-making for optimizing parking infrastructure and enhancing sustainable mobility solutions.
@@
  • ATM: Autoritat de Transport Metropolità
@@ 6.4 Authors’ contribution
AK and VG designed the research project and performed data analysis, model optimization and evaluation. JF and DM performed data analysis, model optimization and evaluation. All authors contributed to the writing and reviewed the manuscript.

### v2

@@ 4.3.1 Fitting the TN Model
For the basic TN model, the objective is to estimate the parameters $\boldsymbol{\theta}=(\mu_{a},\mu_{d},\sigma_{a},\sigma_{d})$ from a dataset consisting of $N$ daily occupancy profiles, $\mathcal{D}~=~\{o_{1:T}\}_{n=1}^{N}$, recorded at $T$ discrete timestamps [^3] In our case, $T=48$ since the dataset has a 30-minute time resolution.. To ensure consistency, all occupancy profiles are normalised such that $\sum_{i=1}^{T}o_{i}=1$.
@@
The optimisation can be carried out by minimising the sum of squares loss
@@ 4.3.2 Fitting the TNL Model
For the TNL model with occupancy limit, the set of fitted parameters $\boldsymbol{\theta}$ also includes a set of threshold parameters $\{\tau_{i}\}_{i=1}^{N}$. Thus, we allow $\tau$ to take a different value for each training day $i$ while keeping the other parameters fixed. This approach ensures that the parameters of the arrival distribution remain independent of the exact moment a parking lot reaches capacity. As a result, the model can not only predict when the capacity limit will be reached but also estimate the number of cars that were unable to find a free parking space. Note, however, that this approach requires a different normalisation of the data. In this case, all occupancy profiles are scaled such that $\max(o_{1:T})=1$. This normalisation is essential for identifying the moments when the parking lots reach full capacity.
@@ 4.4.1 Average Activity Profile
A straightforward method for predicting patterns influenced by circadian rhythms is the use of average activity profiles, as employed in [16]. We adapt this approach by computing the average over normalised, aggregated data, distinguishing between three groups: weekdays (Monday to Thursday), Fridays, and weekends (Saturday and Sunday).
@@ 4.4.2 Linear Regression
The regression model optimises the following objective:
@@ 5.1 Model Fit Quality
We begin by assessing how well the model represents both the training and test data. For this, we use the mean relative proportional error, computed relative to the maximum occupancy of a parking lot. For clarity, we group the test data into three categories based on observed activity patterns (as illustrated in Figure 3): weekdays (Monday to Thursday), Fridays, and weekends (Saturday and Sunday). The goodness of fit of the models in terms of the average loss per training day is given in Tables 5 and 6 in Appendix B.
@@ 5.1.1 Example Fits on Training Data
Figure 7: Example of the TN model fitted to training data for the Vilanova car park. The parameters of the arrival and departure TN distributions are $\mu_{a}=$~06:56h, $\mu_{d}=$~18:40h for weekdays, $\mu_{a}=$~07:02h, $\mu_{d}=$~17:27h for Fridays and $\mu_{a}=$~07:50h for weekends. Note that for the TN model the data is normalised by the area under the curve. Av. Loss indicates average sum of squared loss per day.
@@
Figure 8: Example of the TNL model fitted to training data for the Quatre Camins car park. The corresponding parameters of the arrival and departure distributions are $\mu_{a}=$~07:32h, $\mu_{d}=$~19:25h for weekdays, $\mu_{a}=$~07:43h, $\mu_{d}=$~18:30h for Fridays and $\mu_{a}=$~10:10h and for weekends. Dashed lines indicate stdv of the model curve. The car park usually reaches its capacity limit from Monday to Friday between 8 and 8:30. Note that for the TNL model the data is normalised by the max value.
@@
We first illustrate model fits on the training data using two example car parks. Figure 7 shows the extracted TN model curves for the Vilanova car park, while Figure 8 shows the TNL model curves, incorporating an occupancy limit, for the Quatre Camins car park. The latter reaches maximum capacity on most weekdays and Fridays (since for the weekend profile, this does not happen, the basic TN model is used there).
@@
The dark grey lines represent actual occupancy profiles from the training data. The best-fitting model curves are shown as solid red lines, while blue dashed lines indicate the standard deviation of the model fit. All curves are normalised: in the TN model, by the area under the curve, and in the TNL model, by the maximum value. The latter normalisation is necessary to estimate excess demand in a given car park. For Figure 8, the model curve is plotted using the average excess across all training curves. We observe the best fits for weekday profiles, while Fridays, especially weekends, exhibit greater variability.
@@
The fitted TN model curves and their corresponding parameters are shown in Figure 9. The figure presents model curves for weekdays (left panel) and Fridays (right panel) across all eight car parks, ordered by the difference between the location parameters $\mu$ of the arrival and departure distributions (represented as circle markers). The interval corresponding to $\pm$ one standard deviation $\sigma$ around $\mu$ is indicated by dashed lines and horizontal markers for both the arrival and departure distributions. The exact values of the parameters are given in Appendix B in Table 3 for the TN model and in Table 4 for the TNL model.
@@
A smaller time difference between arrivals and departures is observed on Fridays, primarily due to earlier departure times. This pattern aligns with the common practice of shorter work schedules on Fridays, often without lunch breaks. The most pronounced example is the Prat del Llobregat car park, where the average departure time shifts from approximately 17:20 to 14:20—a three-hour advancement. We exclude the corresponding curves for weekends, which are characterised by more extreme parameter values and greater variability.
@@
Figure 9: Model curves, with parameters for the TN model, for weekends (left panel) and Fridays (right panel) for the different car parks. The dots indicate the mean $\mu$ and the dashed line the corresponding standard deviation interval $\mu-\sigma,\mu+\sigma]$ of the underlying TN distributions for arrival (right) and departure (left).
@@ 5.1.2 Evaluating Model Performance on Test Data
We now analyse the model’s performance by evaluating errors in the test data. Examples of two individual stations are given in Appendix A.
@@
We find that, apart from the four stations that show significant variability on Fridays (Prat Del Llobregat, Cerdanyola, Mollet, and Sant Boi, the latter also on weekends), the average error is always below 4%. The high error on weekends for the Sant Boi car park is caused by party-goers visiting nearby nightlife attractions. This results in a high-occupancy phase during the early morning hours, which is not well captured by the current models. To address this issue, future work could extend the model and use cycles that last longer than 24 hours on weekends and end when the car parks reach their minimum occupancy.
@@ 5.2 Prediction
Next, we analyse the predictive quality of our model. More precisely, at a given moment in time, we aim to predict the evolution of occupancy in the subsequent hours of the day. Again, we measure prediction quality using the relative absolute error with respect to the maximum occupancy of a car park and aggregate weekdays into three groups. We also compare the results with average profiles, which represent the typical activity in a given car park for the corresponding types of days in the training data.
@@
The left panels display real data (dash-dotted lines) along with the predicted occupancy, while the right panels show the corresponding prediction errors for both the model curve (in blue) and the average profile (in orange). The average errors are depicted as dashed lines. The dark grey area represents the data used to generate the prediction.
@@
Figure 11: Three examples of predictions in a car park (Vilanova) with the TN model. Left: The grey areas indicate the known activity used to predict the number of parked cars (red dashed-dotted lines) during the remaining hours of the day. Right: evolution of the relative prediction error in %. Blue lines indicate performance of the TN model, while orange lines show the comparison with an average day-cycle profile.
Figure 12: Three examples of predictions in a car park which fills up (QuatreCamins) with the TNL model. Left: The grey areas indicate the known activity used to predict the number of parked cars (red dashed-dotted lines) during the remaining hours of the day. Right: evolution of the relative prediction error in %. Blue lines indicate performance of the TNL model, while orange lines show the comparison with an average day-cycle profile. Black line gives an estimate of the surplus (how many cars will not or did not fit in the parking).
@@ 5.3 Nowcasting
Table 2: Median nowcasting error in % for the four stations depicted in Figure 14 for weekdays and Fridays comparing the Truncated Normal with Limit (TNL), basic (TN), and Average Profile (Av. Prof.) model with a Linear Regression (L-Reg.). Nowcasting is performed between 7:00 and 15:00 in 30-minute intervals.
@@
6 Policy Recommendations
Based on the findings of this study, we propose the following policy recommendations for transportation planners and municipal authorities involved in managing park-and-ride infrastructure:
  • Use simple, explainable models for decision support: The proposed Truncated Normal (TN) and Truncated Normal with Limit (TNL) models provide accurate occupancy predictions using a minimal set of interpretable parameters. This simplicity facilitates communication with non-technical stakeholders and supports transparent, evidence-based decision-making.
  • Monitor and address unmet demand: The TNL model explicitly estimates the number of vehicles unable to find parking during peak times. Authorities can use this to identify high-demand stations and prioritise infrastructure investments (e.g., expanding parking capacity or improving alternative access) where overflow is consistently observed.
  • Incorporate behavioural insights in planning: Because the model parameters correspond to actual commuter behaviour (e.g., average arrival and departure times), planners can use these insights to design time-specific interventions (such as staggered work hours, variable pricing, or promotional campaigns encouraging off-peak usage) to better balance demand.
  • Enable low-cost scalability using aggregate data: Our approach requires only aggregate occupancy data, which is often easier and cheaper to collect than individual vehicle tracking. This enables scalable and privacy-preserving deployment across a broad set of car parks without major investments in infrastructure.
  • Prioritise regularly saturated Facilities: As shown in our results (e.g., Figure 4), some car parks consistently reach their capacity limits. We recommend incorporating our model outputs into multi-criteria prioritisation frameworks to guide expansions, especially in facilities critical for multimodal commuting corridors.
By aligning parking infrastructure planning with behavioural patterns and unmet demand forecasting, authorities could improve efficiency, user satisfaction, and the overall effectiveness of park-and-ride systems.
@@ 7 Conclusion
Additionally, our work differentiates itself from related studies through its specific focus on the modelling approach. Unlike some studies that prioritise prediction using techniques without an underlying behavioural model [3, 4, 5], our approach is more ambitious, aiming to understand the fundamental behavioural patterns of commuters rather than merely forecasting occupancy trends.
@@
However, we also have to acknowledge two limitations of our study. First, our model assumes that arrival and departure times are independent random processes. This simplification enables tractability and parameter identifiability using aggregate data. While it may not fully capture dependencies present in individual commuter behaviour (e.g., fixed work shifts), the strong regularity observed in occupancy patterns suggests that, at the aggregate level, this assumption remains a reasonable approximation. Future work could explore coupled distributions or joint modelling frameworks if individual-level data becomes available. Secondly, the model is tailored to regular commuting behaviour and may be less suitable in non-routine contexts such as shopping centres or event-driven demand, where more flexible or hybrid models may be required.
Nonetheless, and ultimately, we believe this work will contribute to improved urban planning by offering a practical tool for understanding and predicting commuter behaviour in park-and-ride systems. By balancing simplicity, interpretability, and predictive accuracy, our approach can support data-driven decision-making for optimising parking infrastructure and enhancing sustainable mobility solutions.
@@
  • ATM: Autoritat del Transport Metropolità
@@
7.1 Ethics approval and consent to participate
Not applicable.
7.2 Consent for publication
Not applicable.
@@ 7.6 Authors’ contribution
AK and VG designed the research project and performed data analysis, model optimisation and evaluation. JF and DM performed data analysis, model optimisation and evaluation. All authors contributed to the writing and reviewed the manuscript.
@@ Appendix A Error Analysis for Two Representative Car Parks
Appendix B Model parameters and loss
In this section, we display the values of the parameters obtained when fitting the training data and also the sum of squares loss of the corresponding optimisation processes.
Table 3 lists the parameters for the TN model for all car parks, and Table 4 lists the parameters of the TNL model for the four car parks which reach their capacity limits during weekdays and Fridays. The corresponding values of the average loss functions per day of training data are given in Tables 5 and 6. Note that the values of the average loss function are not directly comparable between the TN and the TNL model, as the two models perform different normalisation strategies.
Table 3: Parameters $\mu_{a}\pm\sigma_{a}$, $\mu_{d}\pm\sigma_{d}$ (in hh:mm units) of the TN Model for the different stations. indicates cases where higher early morning occupancy during weekends (as explained in Section 5.1.2) causes non-interpretable parameter values..
Car Park Weekdays Fridays
arrival departure arrival departure
PratDelLlobregat 08:07 $\pm$00h58 17:21 $\pm$02h50 08:36 $\pm$01h25 14:22 $\pm$03h48
Granollers 07:18 $\pm$00h58 18:21 $\pm$02h01 07:18 $\pm$00h40 16:34 $\pm$01h59
Vilanova 06:56 $\pm$01h16 18:40 $\pm$03h05 07:02 $\pm$01h35 17:27 $\pm$03h33
Mollet 07:02 $\pm$00h49 18:59 $\pm$02h15 07:22 $\pm$01h00 18:29 $\pm$04h39
QuatreCamins 07:17 $\pm$00h43 19:20 $\pm$01h53 07:23 $\pm$00h48 18:19 $\pm$02h57
SantSadurni 07:09 $\pm$01h07 19:23 $\pm$02h22 07:16 $\pm$02h04 18:12 $\pm$03h27
SantBoi 07:29 $\pm$01h03 20:17 $\pm$02h15 07:27 $\pm$01h02 6809:53 $\pm$164h43
Cerdanyola 07:11 $\pm$01h19 21:33 $\pm$05h46 05:52 $\pm$03h23 16:42 $\pm$08h27
Weekends
PratDelLlobregat 08:06 $\pm$03h09 22:55 $\pm$07h54
Granollers 07:26 $\pm$02h34 10786:37 $\pm$249h47
Vilanova 07:50 $\pm$03h00 248917:52 $\pm$730h15
Mollet 08:34 $\pm$02h46 22:39 $\pm$02h13
QuatreCamins 09:29 $\pm$02h08 22:53 $\pm$02h31
SantSadurni 09:49 $\pm$01h58 19:32 $\pm$04h47
SantBoi 00:00 $\pm$07h47 33:11 $\pm$02h09
Cerdanyola 08:29 $\pm$04h08 18:26 $\pm$07h24
Table 4: Parameters $\mu_{a}\pm\sigma_{a}$, $\mu_{d}\pm\sigma_{d}$ (in hh:mm units) and $\tau$ (in %) of the TNL Model for the different stations which reach the occupancy limits. indicates a case where higher early morning occupancy during weekends (as explained in Section 5.1.2) causes non-interpretable parameter values.
Car Park Weekdays Fridays
arrival departure avg. $\tau$ arrival departure avg. $\tau$
Mollet 07:06 $\pm$00h52 19:00 $\pm$02h16 80.66% 07:20 $\pm$00h56 19:28 $\pm$04h07 72.48%
QuatreCamins 07:32 $\pm$00h52 19:25 $\pm$01h51 79.58% 07:43 $\pm$00h55 18:30 $\pm$02h50 73.08%
SantSadurni 07:20 $\pm$01h17 19:20 $\pm$02h30 76.90% 07:08 $\pm$01h47 18:17 $\pm$02h57 66.69%
SantBoi 08:00 $\pm$01h22 20:19 $\pm$02h07 75.75% 11:05 $\pm$02h13 32:07 $\pm$07h27 12.29%
Table 5: Average Loss per day of the TN Model for the different stations.
Car Park Weekdays Fridays Weekends
PratDelLlobregat 0.00179 0.00638 0.00541
Granollers 0.00029 0.00246 0.00346
Vilanova 0.00014 0.00042 0.00174
Mollet 0.00058 0.00269 0.00215
QuatreCamins 0.00026 0.00031 0.00342
SantSadurni 0.00021 0.00089 0.00589
SantBoi 0.00025 0.00057 0.01242
Cerdanyola 0.00209 0.00474 0.01082
Table 6: Average Loss per day of the TNL Model for the different stations.
Car Park Weekdays Fridays
Mollet 0.331 1.072
QuatreCamins 0.141 0.149
SantSadurni 0.135 0.459
SantBoi 0.159 0.395
Figure Legends
Table Legends